Server migration (Resolved)
  • Priority - High
  • Affecting Server - Linux Web #1
  • Hello all,

    We're opening a maintenance window for all web hosting accounts on Linux Web #1 to be migrated to a brand new web server.

    As part of this upgrade there will be a period of unavailability for your site whilst the DNS propagates to the new IP addresses.

    During the upgrade process you'll find that if you're using the MySQL services on the web hosting remotely that they'll cease to respond if you're using the IP address or
    Once the migration is complete you'll either need to make sure the connection information for whatever is connecting to the database uses the domain name as the host, or the new IP address of the web server, which will be

    For those of you interested in the specification of the new and old web server!

    Old server -
    Intel Core i5-4460 3.2Ghz (4C/4T)
    16GB DDR3 1600Mhz RAM
    2x 1TB HDD's in RAID1
    1Gbps network port
    Running CentOS 6.9

    New server -
    Intel Xeon E5-1650v3 3.5Ghz (6C/12T)
    64GB DDR4 2400Mhz ECC RAM
    2x 1TB SSD's in RAID1
    1Gbps network port
    Running CentOS 7.4

    The migration process will begin at 00:30 on Thursday the 2nd of November 2017.

    We expect the migration of all web site data to take several hours till at least 6-7AM on Thursday.

    Once complete we shall update the status of this post to let everyone know. If any problems do arise during the migration process we'll be sure to update this post accordingly, but hopefully everything shall be A-OK!

    ** UPDATE 06:27 AM GMT ** Data migration is still in progress and is only just over 30% done, this could mean that it could be 18:00 GMT before all sites are fully migrated.

    ** UPDATE 03:00 PM GMT ** 
    All data has been sucessfully migrated. Please remember to use the new IP address, or hostname for remote MySQL connections if you are not doing so already. If you're having any other issues please drop in a support ticket on the link below. 

    Kind regards,
    Nitrous Networks Team.

  • Date - 02/11/2017 00:30
  • Last Updated - 02/11/2017 15:01
Bristol DC - Scheduled downtime (Resolved)
  • Priority - High
  • Affecting Other - Affecting services hosted in Bristol, UK
  • Dear customer,

    We're opening a scheduled window for downtime in our Bristol location to carry out some essential maintenance to ensure no unanticipated downtime in the future.

    Start date / time - 11/09/2017 @ 23:59 (BST)
    End date / time - 12/09/2017 @ 00:30 (BST)

    Description of works - We're replacing our fibre WDM unit in Bristol with a new unit due to faults with fans internal to the unit, which if not remedied could cause unanticipated downtime in the future.

    Please note during this time no services hosted within Bristol will be accessible (this should only affect dedicated server / community tier customers located in Bristol, web hosting and UK TS3 servers).

    We shall update this post once the maintenance period has started, and been closed.

    ** 00:27 ** All work has been completed as expected and services are back online.

    Kind regards,
    Nitrous Networks

  • Date - 11/09/2017 23:59 - 12/09/2017 00:30
  • Last Updated - 12/09/2017 00:41
Problems in Atlanta (Resolved)
  • Priority - High
  • Affecting Other - ATL-2
  • Hello,

    A DC technician is currently working on resolving a problem with the CPU within ATL-2 being fixed at 800Mhz and not scaling up to 3.9Ghz like it should be. This was causing significant performance issues for users on this node, and increased latency.

    Update 22:30 BST 10th August


    After swapping this blade with another in the same chassis, it appears this issue is with the entire chassis. We are attempting to get your server moved to another chassis with the hope of resolving the issue.

    Thank You


    Brandon N.

    DC Technician”


    Update at 02:04 AM BST



    I apologize again for all the issues and delay in response. We have resolved the CPU issue with the swap, and the device is in the final stretch of being operable. Our support team is finishing the configurations and should have the device up and ready within the hour.



    Thank You
    Arthur J.
    Data Center Technician


    Update at 02:29 AM BST - Issue is now resolved. Thank you for your patience.

    We’ll update this with more information as soon as we have it.


    Nitrous Team.

  • Date - 10/08/2016 02:00 - 11/08/2016 02:29
  • Last Updated - 11/08/2016 02:28
Scheduled maintenance window (Resolved)
  • Priority - High
  • Affecting Other - Affecting servers hosted in Bristol, England.
  • Dear valued customer,

    This is a scheduled maintenance window notification to inform you of upcoming required downtime in order to upgrade the network infrastructure within our Bristol hosting site.

    Description of work being carried out: Upgrading of aggregate switching stack. Switching from Cisco 3560-X stack, to Juniper EX4200 virtual chassis setup in order to better handle larger volumes of IPv4 / IPv6 traffic.

    Start date & time: 23/05/2016 (DD/MM/YYYY) @ 08:30 AM BST.
    End date & time: 23/05/2016 (DD/MM/YYYY) @ 09:00 AM BST.

    **Update at 10:20PM - The above mentioned upgrades have been completed. Unfortunately it took a little longer than expected because of an issue with the primary file system on one of the RE's but this has been repaired now. Once again thank you for your patience.**


    Thank you for your patience and continued support.

    Kind Regards.
    Nitrous Team.

  • Date - 23/05/2016 08:30 - 23/05/2016 09:00
  • Last Updated - 23/05/2016 10:31
Domains / Hostnames not resolving (Resolved)
  • Priority - High
  • Affecting Server - Linux Web #1
  • Hello,

    We're currently investigating an issue with domain names / hostnames not resolving which is causing issues connecting to websites hosted on this server.

    Upon further inspection the fault appears to be with the domain that we use for our name server addresses.
    This is being looked into by the registrar of the domain, 123-reg and will hopefully be resolved very soon.

    UPDATE 03/11/2015 17:26 GMT - The ticket with our registrar has been escalated, we hope to have a resolution to this fairly soon.

    UPDATE 04/11/2015 12:33 GMT - The name servers for the domain were reset yesterday. We were unable to change them without encountering an error to point back to our servers. 123-reg have escalated the issue further to their development team. Apologies for the delay.

    UPDATE 04/11/2015 14:34 GMT - Name servers have been set correctly now by the registar. They're still looking into the root cause of why it happened in the first place. Your website / hostname should be working now, or will be working for you in the next few hours.

    Apologies for any inconvenience caused by this.

    Nitrous Team.

  • Date - 03/11/2015 16:00
  • Last Updated - 04/11/2015 15:14
Server down (Resolved)
  • Priority - Critical
  • Affecting Server - Linux Web #1
  • Hello,

    We're currently investigating an issue with our web server Linux Web #1 being offline.

    We'll keep you updated by posting updates here.

    31/10/2015 00:09 GMT - RAID was found to be degraded. Faulty drive has been replaced and the mirror is re-building.
    Please allow for up to 24 hours for this to complete. During which time sites will be accessible but may be substantially slower than normal.

    Kind Regards,
    Nitrous Networks.

  • Date - 30/10/2015 22:54 - 31/10/2015 17:19
  • Last Updated - 31/10/2015 00:12
Scheduled Maintenance (Resolved)
  • Priority - Critical
  • Affecting Other - Customers hosted within RACK6-7
  • Hello,

    Due to our company's rapid ongoing expansion and upgrades within our Bristol data centre we need to schedule in a short downtime window in order to migrate 2 racks of servers into a newly constructed server aisle as part of refurbishment plans in our data hall. Please note, only people that have been emailed or have an alert on their control panel will be affected by this.


    Start time (DD/MM/YY HH:MM): Sunday16/08/2015 23:30 PM

    End Time (DD/MM/YY HH:MM): Monday 17/08/2015 00:30 PM

    **This was all completed by 01:30AM GMT without any problems, thank you for your patience**


    Reason for downtime: Relocation of servers from 2 racks within older section of data hall in order to support refurbishment plans over the next month.

    Please note servers will start to be powered down from 23:30PM on wards and should start coming back online anytime from 00:01 on wards.

    Thank you for your patience.

  • Date - 16/08/2015 23:30 - 17/08/2015 00:45
  • Last Updated - 17/08/2015 14:11
Network maintenance (Resolved)
  • Priority - High
  • Affecting Other - All UK Servers
  • Dear Customer,

    Please accept this as notification for an emergency maintenance window to be conducted at the following time:

    Date: 8th December, 2014
    Time: 00:45
    Duration: 10 minutes

    During the above window and on the advice of Juniper TAC, we will be replacing a failing line card on our Juniper MX located in Telehouse North. This is a proactive maintenance window to ensure that the line card does not expectedly fail in operation.


    Nitrous Team.

  • Date - 08/12/2014 01:00 - 08/12/2014 01:10
  • Last Updated - 08/12/2014 01:44
Network maintenance (Resolved)
  • Priority - Critical
  • Affecting Other - All UK Servers
  • Hello everyone,

    We have planned scheduled maintenance set for today at our Bristol location affecting all servers hosted on the 185.38.148.XXX and 185.38.149.XXX IP range.

    Starting at: 05/12/2014 @ 09:00
    Ending by: 05/12/2014 @ 10:30

    Update at 11:35 05/12/2014 - Engineers were caught in mid morning traffic so ended up getting to the site around 45 minutes later than normal. They're just currently finishing up now and are just running some final tests.

    what’s going on? The reason behind this maintenance is to finalize the fail over point of our redundant fibre link that runs a different route to our current national fibre that runs from our facility in Bristol back to the docklands in London. Optic engineers will be on site from just before 09:00 today and will be testing various points and simulating loss of light which would happen if the fibre was cut on our primary feed, in order to test the auto fail-over onto the secondary feed that runs from Bristol via Newport, and then back to London into our routing equipment at Telehouse.

    We're told this shouldn't take longer than an hour but we've scheduled a 90 minute window to give a bit of head room.

    Sorry for any inconvenience this may cause you but this is necessary to provide an improved quality of service to avoid what has happened in the past when fibre cuts have occurred.

    Thanks again and stay awesome!

    Nitrous Networks Team.

  • Date - 05/12/2014 09:00 - 05/12/2014 10:30
  • Last Updated - 05/12/2014 13:14
Network Downtime (Resolved)
  • Priority - Medium
  • Affecting Other - Scheduled maintenance window to NextConnex fibre (UK location)
  • **Please note this only applies to customers hosted within the UK at our Bristol location**

    **Update @ 2014-10-16 02:09:33 - Our fibre has been spliced and re-connected and BGP sessions are coming back up. You should still consider this at risk until 7AM GMT once the all clear has been issued. Thank you again for your patience **

    Hello everyone,

    Following up from the outage on the 2nd of October with the fibre cut on our fibre carriers network they've opened a short notice scheduled maintenance window with us for the 16th of October. Please see below for the full message sent from them:-

    "Dear NextConnex Customer,

    Please see below the details of emergency works.

    Short Notice Full outage Maintenance Window.
    Start Date & Time 16-10-2014 00:01
    End Date & Time 16-10-2014 07:00

    Service Affecting?: Yes

    Description of Work:
    This is a CAT D FULL OUTAGE Planned Works Notification. Due to the proposed Bridge deck replacement at Church Bridge, Church road, Frampton Cotterell, the primary duct has been temporarily re-routed in order for the bridge deck replacement works to be undertaken. An interruption cable is in place. These works are to re-route the cable back through the new bridge deck & recover the interruption cable. It is planned that the service outage will begin at 00.01hrs on 16/10/2014 when the new cable will be integrated into the mid-span joint closures. The splicers will then work to a splicing plan making sure all tubes containing customer fibres are spliced back as a priority. Each tube will be fully spliced and returned to the splice enclosure/tray before commencing work on the next tube. All fibres containing customer circuits will be restored by 07:00hrs, if there are any unforeseen circumstances then the Geo NOC will advise the customer prior to 07.00hrs on 16/10/14. Total cabling & splicing time is expected to take 7 hours with final completion estimated to be 07:00hrs on 16/10/14

    NextConnex apologizes for any inconvenience caused by these works.


    NextConnex NOC"

    During which time the network to our Bristol site will be unavailable whilst they're working on the fibre.

    We're currently working through some plans with NextConnex for a purposed fail over / redudent routes in to our Bristol facility back to the docklands.

    Update 15/10/2014 - We've had a meeting with NextConnex today and we've come up with a plan to split the wavelegnth and take a seperate signal to NGD, a data centre in Newport and pick up transit from another carrier there for fail over. This will ensure redundecy on both a L1 carrier and IP transit level as the redundent feed would also be part of a national MPLS network with multiple routes to itself. We will be going over this with NextConnex and hopefully deploying this solution within the next 6-8 weeks depending on availability for optic engineers etc.

    Thank you for your continued patience and support.

    Nitrous Networks.

  • Date - 16/10/2014 00:01 - 16/10/2014 13:07
  • Last Updated - 16/10/2014 02:58
Fibre Cut (Resolved)
  • Priority - Critical
  • Affecting System - Fibre cut on NextConnex network
  • Hello everyone,

    Sorry for the delay in sending this out, we've been busy trying to crack down on support tickets and keeping in contact with our fibre carrier.

    As many of you will be aware by looking at our Twitter that our fibre carrier that provides our route from Bristol to the docklands in London is currently experiencing a fibre cut between their Bristol and Swindon PoP's which has interrupted traffic to and from our Bristol location.

    The best thing to do at this point in time is follow any updates as they come in on our Twitter account here:-

    Currently we suspect the ETA for repair to be within 18:30 and 19:30. Due to current rush hour traffic conditions the first team of field engineers is only just arriving to the location.

    Thank you for your patience.

    *Issue has now been resolved**

    Nitrous Networks.

  • Date - 02/10/2014 14:30 - 02/10/2014 23:47
  • Last Updated - 02/10/2014 23:47